# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v1

@@ II.1 Cyclic representations in different metric spaces
Consider two faithful cyclic representations of the system $(\mathcal{A},\omega)$: (¯$(\bar{\mathcal{H}}_{\omega},\bar{\pi}_{\omega},\bar{\Phi}_{\omega})$ and $(\mathcal{H}_{\omega},\pi_{\omega},\Phi_{\omega})$. Let us assume that the two Hilbert spaces are different metric spaces where $\mathcal{H}_{\omega}$ is endowed with metric $G$ and ¯$\bar{\mathcal{H}}_{\omega}$ endowed with metric $G^{\prime}$. Therefore $\mathcal{H}_{\omega}=\mathcal{H}_{G}$ and ¯$\bar{\mathcal{H}}_{\omega}=\mathcal{H}_{G^{\prime}}$. Since these are cyclic representations, they must be unitarily equivalent, hence there exists a unitary map [^1] By “unitary” we mean an invertible isometry. In the most general sense dual of a unitary is equal to its inverse $T:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$ between the two Hilbert spaces whose action is given by $\Psi_{A}\rightarrow T\Psi_{A}$, where $T$ is a bounded and positive operator. Therefore we should have $\langle\psi_{A}|\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}=\langle T\psi_{A}|T\psi_{A}\rangle_{\mathcal{H}_{G}}=\langle\psi_{A}|T^{*}T\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}$, or in other words $T^{*}T=1$. One can check that the dual operator $T^{*}$ is given by $T^{*}=G^{\prime-1}T^{\dagger}G$. Combining this with the unitarity property, we have the relation $G^{\prime}=T^{\dagger}GT$. We had discussed this relation in the introduction of the metric formalism. Here we have seen that this is a consequence of the the unitary equivalence of the cyclic representation of the same underlying algebra. The operator $T$ intertwines the algebras $\pi_{G}(\mathcal{A})$ and $\bar{\pi}_{G^{\prime}}(\mathcal{A})$: $\pi_{G}(A)=T\pi_{G^{\prime}}(A)T^{-1}\,\forall\,A\in\,\mathcal{A}$. Note that the Hermitisization operation $\eta:\mathcal{H}_{G}\rightarrow\mathcal{H}:\psi\rightarrow\eta\psi$ is a special case of such a unitary operation. Therefore, given a system $(\mathcal{A},\omega)$, it can have unitarily equivalent faithful cyclic representations in different metric Hilbert spaces and always has a trivial metric (Hermitian) description  (see also  [27]).
@@ II.2 Partitions of a system
Let us now restrict ourselves to defining subsystems in representations $(\mathcal{H}_{G},\pi_{G})$, where $G\neq G_{1}\otimes G_{2}$, while we do have $\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$, i.e. the bi-partition in the non-trivial metric Hilbert space is not a genuine bi-partition. An observable $\pi_{G}(A)$ under this representation, is also an observable in the Hilbert space $\mathcal{H}_{G^{\prime}}$ if $\pi_{G}(A)$ commutes with the unitary map $T:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$. Therefore, given an observable, there can be an infinite number of metric spaces to choose from. However, if we fix a certain set of observables $\{\pi_{G}(A_{i})\}_{i}$ to probe the system, such that the only operator that commutes with all the elements of the set is a multiple of identity, then the metric space under which this set remains a set of observables is fixed  [29]. Such a set is called an irreducible set of operators as they do not have any common invariant subspace. This discussion points to the relationship between the choice of the metric operator and the choice of observables. The question then naturally arises, given a partitioning of an algebra, to what extent is the metric operator fixed? The following theorem answers this question.
@@
The partitioning of a system is fixed by a choice of an equivalence class of metric operator.
@@
We will show this for a bi-partition of an algebra $\mathcal{A}$. Let $\mathcal{A}_{1}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathcal{O}_{A}\otimes\mathbb{1}\}$ and $\mathcal{A}_{1}^{\prime}=\{A\in\mathcal{A}:\eta^{\prime}\pi_{G^{\prime}}(A)\eta^{\prime-1}=\mathcal{O}_{A}^{\prime}\otimes\mathbb{1}\}$ be the two sub-algebras defined w.r.t. two different metric operators $G=\eta^{2}$ and $G^{\prime}={\eta^{\prime}}^{2}$ respectively. One defines the complementary sub-algebras $\mathcal{A}_{2}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{A}\}$ and similarly, $\mathcal{A}_{2}^{\prime}$. This gives us two bi-partitions of the algebra $\mathcal{A}$. Now we will determine the condition for the equality of the two algebras $\mathcal{A}_{1}\sim\mathcal{A}_{1}^{\prime}$. Let $X\in\mathcal{A}_{1}$, so that $\eta\pi_{G}(X)\eta^{-1}=\mathcal{O}_{X}\otimes\mathbb{1}$. Let $T:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$ be an invertible isometry, then we know $\pi_{G^{\prime}}(X)=T^{-1}\pi_{G}(X)T$ and $\eta^{\prime}=U\eta T$. Combining these two relations we have $\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=U\eta\pi_{G}(X)\eta^{-1}U^{\dagger}$. Therefore $X\in\mathcal{A}_{1}^{\prime}$, or in other words $\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=\mathcal{O}_{X}^{\prime}\otimes\mathbb{1}$, if and only if $U=U_{A}\otimes U_{B}$. We define the equivalence class of metric operator equivalent class of metric operators
$[G]\equiv\{G^{\prime}=\eta^{\prime 2}=T^{\dagger}GT;\,\eta^{\prime}=(U_{A}\otimes U_{B})\eta T\,\text{where }G=\eta^{2}\}.$ (1)
Therefore $\mathcal{A}_{1}\subset\mathcal{A}_{1}^{\prime}$ if and only if the corresponding metric operators belong to the same equivalence class. Proof for $\mathcal{A}_{1}^{\prime}\subset\mathcal{A}_{1}$ is exactly the same. Thus a bi-partition of an algebra is uniquely fixed by an equivalence class of metric operator. ∎
@@
As a consequence of this theorem, we see that while the choice of the metric does not influence the system’s properties, viz. the expectation values of the observables, it can change the properties of the subsystem, e.g. the entanglement between the subsystems. This is not unlike in the case of global unitary rotation on the system, which is simply a choice of a global reference frame, which does not change the expectation values on the system, but does change the subsystems’ properties like the entanglement between them. On the other hand, a local unitary transformation keeps the subsystem properties intact. The above theorem explains this observation from an algebraic point of view: performing a local unitary transformation in the trivial metric Hilbert space is the same as choosing a metric space from an equivalence class of metric operators, and therefore keeps the underlying algebraic partitioning the same.
@@
Tensor product structure on $\mathcal{H}_{G}$: A tensor product structure (TPS)  [17] on a Hilbert space $\mathcal{H}_{G}$ is an isomorphism $\varphi:\mathcal{H}_{G}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ where $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{2})=\dim(\mathcal{H}_{G})$. We know that this isomorphism given by the map  $\varphi:\,|\psi\rangle=\eta|\psi\rangle$. Consider another Hilbert space $\mathcal{H}_{G^{\prime}}$ with the metric $G^{\prime}$ and a TPS $\varphi^{\prime}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}:\psi\rightarrow\eta^{\prime}\psi$. There is an isometry between the two Hilbert spaces, as defined in previous sections, is $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:\,\psi\rightarrow T|\psi\rangle$ where the invertible operator $T$ satisfies $T^{\dagger}GT=G^{\prime}$. Therefore, the map $\varphi^{\prime}\circ\mathcal{T}^{-1}:\psi\rightarrow\eta^{\prime}T^{-1}\psi$ defines another TPS over the Hilbert space $\mathcal{H}_{G}$. Since all these maps are isometries, one can show that $\eta^{\prime}=U\eta T$ for a given unitary operator $U$. This gives us a family of TPS on $\mathcal{H}_{G}$ parametrized by the operator $\eta^{\prime}$, or equivalently by $U$. If the unitaries parameterizing the TPS are local transformations, i.e. of the form $U=U_{1}\otimes U_{2}$, we say they generate equivalent TPS [^2] Following the terminology of  [17]. From theorem 1 we see that:
@@
Remark: a family of equivalent TPS on $\mathcal{H}_{G}$ corresponds to the equivalence class of metric operators $[G]$ and consequently the same partitioning of the total algebra.
@@ II.3 Partial trace
The entropy of the sub-system An important result in C-* algebra then says that a cyclic representation of $(\mathcal{A},\omega)$ is irreducible if and only if the state $\omega$ is a pure state over $\mathcal{A}$. The representation of a mixed state $(\mathcal{A},\omega)$ is therefore reducible into invariant sub-spaces $\mathcal{H}_{\omega}^{k}$ : $\pi_{\omega}=\oplus_{k}\ \pi_{\omega}^{k}$, and the Hilbert space decomposes into the invariant sub-spaces $\mathcal{H}_{\omega}=\oplus_{k}\ \mathcal{H}_{\omega}^{k}$. This decomposition can be used to construct a spectral decomposition of the state, where the invariant sub-spaces are the eigen-sub-spaces of the state: $\rho_{\omega}=\oplus e_{k}\rho_{k}$. Here $e_{k}$ are the eigenvalues of the density matrix, and if the eigen-spectrum has no degeneracy, then $\rho_{k}$ are pure states. This corresponds to the preparation of the state $\omega=\sum e_{k}\omega_{k}$. The entropy of this preparation, which we define as the Shanon entropy of the vector $\vec{e}$: $H(\lambda)=-\sum{e_{k}\ln e_{k}}$ is therefore equal to the von-Newmann entropy of the corresponding density matrix $S_{VN}=\rho_{\omega}\ln(\rho_{\omega})$. Complications that arise with the degeneracies in the spectrum have been discussed in the appendix C (see also [33, 34]).
Consider a special case, as discussed in  [30], where the subsystem of interest is the invariant subspace of the representation. Given a Hilbert space of the form $\mathcal{H}=\oplus_{\alpha}\ \mathcal{H}_{\alpha}$ with the metric $G=\oplus_{\alpha}\ G_{\alpha}$, consider a state $\rho_{\omega}=\oplus_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$. If all $G_{\alpha}$ operators were the same, a partial trace can be given by the density metric $\rho_{\omega}^{r}=\sum_{\alpha}\lambda_{\alpha}\rho_{\omega}^{\alpha}\,\in\mathcal{H}_{\alpha}$. However, such a summation of density matrices is not well defined when they belong to different metric spaces. In that one can define the reduced state $\rho^{r}(\beta)=\sum_{\alpha}\lambda_{\alpha}T_{\alpha,\beta}\ {\rho_{\omega}}^{\alpha}{T_{\alpha,\beta}}^{-1}$ on the Hilbert space $\mathcal{H}_{\beta}$. This defines a unique CPTP map between $\mathcal{B}(\oplus\mathcal{H}_{\alpha})\rightarrow\mathcal{B}(\mathcal{H}_{\beta})$. Here, $T_{\alpha,\beta}$ is an isometric operation between the Hilbert spaces $\mathcal{H}_{\alpha}$ and $\mathcal{H}_{\beta}$ and can be written as $T=U\eta_{\alpha}^{-1}\eta_{\beta}$, where $U$ is a unitary. Of course, this description of the reduced state is compatible with the partial trace by Hermitisization in the sense that: $\eta_{\beta}\rho^{r}(\beta)\eta^{-1}_{\beta}=tr_{X}(\eta\rho_{\omega}\eta^{-1})$.

### v2

@@ II.1 Cyclic representations in different metric spaces
Consider two faithful cyclic representations of the system $(\mathcal{A},\omega)$: (¯$(\bar{\mathcal{H}}_{\omega},\bar{\pi}_{\omega},\bar{\Phi}_{\omega})$ and $(\mathcal{H}_{\omega},\pi_{\omega},\Phi_{\omega})$. Let us assume that the two Hilbert spaces are different metric spaces where $\mathcal{H}_{\omega}$ is endowed with metric $G$ and ¯$\bar{\mathcal{H}}_{\omega}$ endowed with metric $G^{\prime}$. Therefore $\mathcal{H}_{\omega}=\mathcal{H}_{G}$ and ¯$\bar{\mathcal{H}}_{\omega}=\mathcal{H}_{G^{\prime}}$. Since these are cyclic representations, they must be unitarily equivalent, hence there exists a unitary map [^1] By “unitary” we mean an invertible isometry defined everywhere on the Hilbert space. In the most general sense dual of a unitary is equal to its inverse $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$ between the two Hilbert spaces whose action is given by $\Psi_{A}\rightarrow T_{U}\Psi_{A}$, where $T_{U}$ is a bounded and positive operator. Therefore we should have $\langle\psi_{A}|\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}=\langle T_{U}\psi_{A}|T_{U}\psi_{A}\rangle_{\mathcal{H}_{G}}=\langle\psi_{A}|T_{U}^{*}T_{U}\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}$, or in other words $T_{U}^{*}T_{U}=1$. One can check that the dual operator $T_{U}^{*}$ is given by the relation $T_{U}^{*}=G^{\prime-1}T^{\dagger}G$. Combining this with the unitarity property, we have the relation $G^{\prime}=T_{U}^{\dagger}GT_{U}$. We had discussed this relation in the introduction of the metric formalism. Here we have seen that this is a consequence of the the unitary equivalence of the cyclic representation of the same underlying algebra. The operator $T_{U}$ intertwines the algebras $\pi_{G}(\mathcal{A})$ and $\pi_{G^{\prime}}(\mathcal{A})$ such that $\pi_{G}(A)=T_{U}\pi_{G^{\prime}}(A)T_{U}^{-1}\,\forall\,A\in\,\mathcal{A}$. Note that the Hermitisization operation $\eta:\mathcal{H}_{G}\rightarrow\mathcal{H}:\psi\rightarrow\eta\psi$ is a special case of such a unitary operation. Therefore, given a system $(\mathcal{A},\omega)$, it can have unitarily equivalent faithful cyclic representations in different metric Hilbert spaces and always has a trivial metric (Hermitian) description  (see also  [28]).
@@ II.2 Partitions of a system
Note that the $C^{*}$-subalgebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ can be replaced by the commuting observable algebras $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. These are the so-called Jordan-Lie-Banach algebras, whose complexification gives us the corresponding $C^{*}$-algebra  [30].
@@
Let us now restrict ourselves to defining subsystems in representations $(\mathcal{H}_{G},\pi_{G})$, where $G\neq G_{1}\otimes G_{2}$, while we do have $\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$, i.e. the bi-partition in the non-trivial metric Hilbert space is not a genuine bi-partition. An observable $\pi_{G}(A)$ under this representation, is also an observable in the Hilbert space $\mathcal{H}_{G^{\prime}}$ if $\pi_{G}(A)$ commutes with the unitary map $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$. Therefore, given an observable, there can be an infinite number of metric spaces to choose from. However, if we fix a certain set of observables $\{\pi_{G}(A_{i})\}_{i}$ to probe the system, such that the only operator that commutes with all the elements of the set is a multiple of identity, then the metric space under which this set remains a set of observables is fixed  [31]. Such a set is called an irreducible set of operators as they do not have any common invariant subspace. This discussion points to the relationship between the choice of the metric operator and the choice of observables. The question then naturally arises, given a partitioning of an algebra, to what extent is the metric operator fixed? The following theorem answers this question.
@@
For a system $\{\mathcal{A},\omega\}$, any two cyclic representations $\pi$ and $\pi^{\prime}$, with representation spaces $\mathcal{H}_{G}$ and $\mathcal{H}_{G^{\prime}}$ respectively, correspond to the same bi-partitioning into system 1 and 2 if and only if they belong to the equivalence class given by:
$[\pi]\equiv\{\pi^{\prime}(\cdot)=T_{U}^{-1}\pi(\cdot)T_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta\}.$ (1)
Here $U_{1}$ and $U_{2}$ are unitary operators corresponding to the two subsystems.
@@
Let $\mathcal{A}_{\pi}^{1}=\{A\in\mathcal{A}:\eta\pi(A)\eta^{-1}=\mathcal{O}_{A}\otimes\mathbb{1}\}$ and $\mathcal{A}_{\pi^{\prime}}^{1}=\{A\in\mathcal{A}:\eta^{\prime}\pi_{G^{\prime}}(A)\eta^{\prime-1}=\mathcal{O}_{A}^{\prime}\otimes\mathbb{1}\}$ be the two sub-algebras defined w.r.t. two different metric operators $G=\eta^{2}$ and $G^{\prime}={\eta^{\prime}}^{2}$ respectively. One defines the complementary sub-algebras $\mathcal{A}_{\pi}^{2}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{A}\}$ and similarly, $\mathcal{A}_{\pi^{\prime}}^{2}$. This gives us two bi-partitions of the algebra $\mathcal{A}$ into $\mathcal{A}^{1}_{\pi}\vee\mathcal{A}^{2}_{\pi}$ and $\mathcal{A}^{1}_{\pi^{\prime}}\vee\mathcal{A}^{2}_{\pi^{\prime}}$.
Now we will determine the condition for the equality of the two algebras $\mathcal{A}_{\pi}^{1}\sim\mathcal{A}_{\pi^{\prime}}^{1}$. Let $X\in\mathcal{A}_{\pi}^{1}$. We know that the operators under the two Hilbert spaces are related by $\pi^{\prime}(X)=T_{U}^{-1}\pi(X)T_{U}$ such that $\eta^{\prime}T_{U}^{-1}=U\eta$. Combining these two relations we have $\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=U\eta\pi_{G}(X)\eta^{-1}U^{\dagger}$. Therefore $X\in\mathcal{A}_{1}^{\prime}$ if and only if $U\mathcal{O}_{X}\otimes\mathbb{1}U^{\dagger}=\mathcal{O}_{X}^{\prime}\otimes\mathbb{1}$ for the given unitary $U$. Similarly we have $\mathcal{A}_{\pi}^{2}\sim\mathcal{A}_{\pi^{\prime}}^{2}$ if and only if $U(\mathbb{1}\otimes\mathcal{O}_{Y})U^{\dagger}=\mathbb{1}\otimes\mathcal{O}_{Y}^{\prime}$ for all $Y\in\mathcal{A}^{2}_{\pi^{\prime}}$. Unitaries satisfying these conditions must always be of the form $U=U_{1}\otimes U_{2}$ (see e.g. [32, 33]). Therefore, the two partitions induced by the representations $\pi$ and $\pi^{\prime}$ are equivalent if and only if the interwining isometry $T_{U}$ satisfies the relation $\eta^{\prime}T_{U}=U_{1}\otimes U_{2}\eta$. It is straightforward to see that representations related by such an intertwiner form an equivalence class. ∎
In a typical experiment, the pseudo-Hermitian Hamiltonian is simulated through post-selection or equal gain-loss mechanisms. The form of the Hamiltonian is fixed by the experimentalist, and therefore one is restricted to a class of metric spaces compatible with this Hamiltonian. Such metric spaces are fixed by the isometry $T_{U}$, or equivalently, the unitary $U$. As a simple corollary of the above theorem, the partitioning of the system is then fixed by the choice of the metric space.
Corollary 1.1.
Given a pseudo-Hermitian operator $H$ such that $H^{\dagger}G=GH$, the choices of the compatible representation spaces $\mathcal{H}_{G^{\prime}}$ with equivalent partitioning are those belonging to the equivalence class [G] defined by
$[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta\,\text{ and }\ [T_{U},H]=0\}.$ (2)
Proof.
Given two cyclic representations $(\mathcal{H}_{G},\pi)$ and $(\mathcal{H}_{G^{\prime}},\pi^{\prime})$, such that $H=\pi(X)=\pi^{\prime}(X)$, therefore the representation spaces are $H$-compatible. There exists an isometry $T_{U}$ between the Hilbert spaces such that $\pi=T_{U}\pi^{\prime}T_{U}^{-1}$ which implies $[\pi(X),T_{U}]=0$. From the previous theorem, the two sub-algebra decompositions $\mathcal{A}_{\pi}^{1}\vee\mathcal{A}_{\pi}^{2}$ and $\mathcal{A}_{\pi^{\prime}}^{1}\vee\mathcal{A}_{\pi^{\prime}}^{2}$ are equivalent if and only if $\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta$. ∎
@@
As a consequence of the above results, we see that while the choice of the metric does not influence the system’s properties, viz. the expectation values of the observables, it can change the properties of the subsystem, e.g. the entanglement between the subsystems. This is not unlike in the case of global unitary rotation on the system, which is simply a choice of a global reference frame, which does not change the expectation values on the system, but does change the subsystems’ properties like the entanglement between them. On the other hand, a local unitary transformation keeps the subsystem properties intact. The above theorem explains this observation from an algebraic point of view: performing a local unitary transformation in the trivial metric Hilbert space is the same as choosing a metric space from an equivalence class of metric operators, and therefore keeps the underlying algebraic partitioning the same.
@@
Tensor product structure on $\mathcal{H}_{G}$: A tensor product structure (TPS)  [17] on a Hilbert space $\mathcal{H}_{G}$ is an isomorphism $\varphi:\mathcal{H}_{G}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ where $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{2})=\dim(\mathcal{H}_{G})$. We know that this isomorphism given by the map  $\varphi:\,|\psi\rangle=\eta|\psi\rangle$. Consider another Hilbert space $\mathcal{H}_{G^{\prime}}$ with the metric $G^{\prime}$ and a TPS $\varphi^{\prime}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}:\psi\rightarrow\eta^{\prime}\psi$. We have also seen that there is an isometry between the two Hilbert spaces, as defined in previous sections, is $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:\,\psi\rightarrow T_{U}|\psi\rangle$. Therefore, the map $\varphi^{\prime}\circ\mathcal{T}^{-1}:\psi\rightarrow\eta^{\prime}T_{U}^{-1}\psi$ defines another TPS over the Hilbert space $\mathcal{H}_{G}$. This gives us a family of TPS on $\mathcal{H}_{G}$ parametrized by the operator $\eta^{\prime}$, or equivalently by $U$. If the unitaries parameterizing the TPS are local transformations, i.e. of the form $U=U_{1}\otimes U_{2}$, we say they generate equivalent TPS [^2] Following the terminology of  [17]. From corollary 1.1 we see that:
@@
Remark: a family of equivalent TPS on $\mathcal{H}_{G}$ corresponds to the equivalence class of metric operators $[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{A}\otimes U_{B})\eta\},$ and consequently the same partitioning of the total algebra.
@@ II.3 Partial trace
The entropy of the sub-system An important result in C-* algebra then says that a cyclic representation of $(\mathcal{A},\omega)$ is irreducible if and only if the state $\omega$ is a pure state over $\mathcal{A}$. The representation of a mixed state $(\mathcal{A},\omega)$ is therefore reducible into invariant sub-spaces $\mathcal{H}_{\omega}^{k}$ : $\pi_{\omega}=\oplus_{k}\ \pi_{\omega}^{k}$, and the Hilbert space decomposes into the invariant sub-spaces $\mathcal{H}_{\omega}=\oplus_{k}\ \mathcal{H}_{\omega}^{k}$. This decomposition can be used to construct a spectral decomposition of the state, where the invariant sub-spaces are the eigen-sub-spaces of the state: $\rho_{\omega}=\oplus e_{k}\rho_{k}$. Here $e_{k}$ are the eigenvalues of the density matrix, and if the eigen-spectrum has no degeneracy, then $\rho_{k}$ are pure states. This corresponds to the preparation of the state $\omega=\sum e_{k}\omega_{k}$. The entropy of this preparation, which we define as the Shanon entropy of the vector $\vec{e}$: $H(\vec{e})=-\sum{e_{k}\ln e_{k}}$ is therefore equal to the von-Newmann entropy of the corresponding density matrix $S_{VN}=\rho_{\omega}\ln(\rho_{\omega})$. Complications that arise with the degeneracies in the spectrum have been discussed in the appendix B (see also [34, 35, felski]).
Consider a special case, as discussed in  [37], where the subsystem of interest is the invariant subspace of the representation. Given a Hilbert space of the form $\mathcal{H}=\oplus_{\alpha}\ \mathcal{H}_{\alpha}$ with the metric $G=\oplus_{\alpha}\ G_{\alpha}$, consider a state $\rho_{\omega}=\oplus_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$. If all $G_{\alpha}$ operators were the same, a partial trace can be given by the density metric $\rho_{\omega}^{r}=\sum_{\alpha}\lambda_{\alpha}\rho_{\omega}^{\alpha}\,\in\mathcal{H}_{\alpha}$. However, such a summation of density matrices is not well defined when they belong to different metric spaces. In fact, one is required to perform a parallel transport of the operators to another metric space in order to add them. The isometry $T_{U}$ defined earlier provides such a metric-compatible parallel transport of operators. For this example, one can define the reduced state $\rho^{r}(\beta)=\sum_{\alpha}\lambda_{\alpha}T_{\alpha,\beta}\ {\rho_{\omega}}^{\alpha}{T_{\alpha,\beta}}^{-1}$ on the Hilbert space $\mathcal{H}_{\beta}$. Here, $T_{\alpha,\beta}$ is an isometric operation between the Hilbert spaces $\mathcal{H}_{\alpha}\to\mathcal{H}_{\beta}$ and can be written as $T_{\alpha,\beta}=\eta_{\beta}^{-1}U\eta_{\alpha}$ for a unitary operator $U$. This procedure defines a unique CPTP map between $\mathcal{B}(\oplus\mathcal{H}_{\alpha})\rightarrow\mathcal{B}(\mathcal{H}_{\beta})$ and is a consistent definition of a partial trace. Note that the reduced state thus defined is independent of the unitary operators $U$ or equivalently the choice of isometries $T_{\alpha,\beta}$. Indeed, if one chooses $U=\mathbb{1}$ for all $T_{\alpha,\beta}$, the partial trace is compatible with the partial trace by Hermitisization: $\eta_{\beta}\rho^{r}(\beta)\eta^{-1}_{\beta}=tr_{X}(\eta\rho_{\omega}\eta^{-1})$.